Write an equation to solve this: If my age is 44 and my sisters age is 48 what age will we each be when our total combined age is 150?
Answer by Maths68(1474)

Write an equation to solve this: If my age is 44 and my sisters age is 48 what age will we each be when our total combined age is 150?
Let
Total Combined age will be 150 after = x years

After x years
Your age = 44+x
Your sister's age = 48+x

Sum of ages
44+x+48+x=150
2x+92=150
2x=150-92
2x=58
2x/2=58/2
x=29
After 29 years total combined age will be 150
Your age = 44+x = 44+29 = 73
Your sister's age = 48+x = 48+29 = 77
Your age + Your sisters age = 150
73+77=150
150=150